OpenStudy (anonymous):

A restaurant has 10 chicken dishes on the menu. It has 7 chicken dishes with pasta, and 15 pasta dishes. What is the probability that a dish has pasta given that it is chicken?

OpenStudy (anonymous):

10 + 7 + 15 = 32 chicken : 10/32

OpenStudy (anonymous):

you can simplify

OpenStudy (anonymous):

5/16

OpenStudy (anonymous):

Did that help ? @KierseyClemons

OpenStudy (anonymous):

it's a decimal lol

OpenStudy (anonymous):

so change 5/16 into a decimal by simply dividing it

OpenStudy (anonymous):

0.3125

OpenStudy (anonymous):

the choices

OpenStudy (anonymous):

My bad i read the question wrong it says what is the probability that a dish has PASTA given that it is CHICKEN

OpenStudy (anonymous):

Not what is the probability that a dish is chicken

OpenStudy (paki):

it would be 15/32= 0.4.... @Shay17 what you say about this...?

OpenStudy (anonymous):

I say it's actually C. 0.7

OpenStudy (paki):

how...?

OpenStudy (anonymous):

There are 22 pasta dishes altogether 15 regular pasta dishes and 7 chicken with pasta dishes

OpenStudy (paki):

@shereenkhan what you say about this question.... will you help us...?

OpenStudy (anonymous):

actually i am counfuse.

OpenStudy (anonymous):

i am trying.

OpenStudy (paki):

@Shay17 then it would be 22/32 = 0.6...

OpenStudy (anonymous):

15/29+ 0.5

OpenStudy (anonymous):

@KierseyClemons choose C I promise it's the right answer :)

OpenStudy (anonymous):

@paki It's actually 0.6875 which is 0.70

OpenStudy (paki):

good.... yup, now you are right..... :-) @Shay17

OpenStudy (anonymous):

@KierseyClemons I hope you didn't pick the wrong answer sorry if I confused you at first

OpenStudy (paki):

@shereenkhan kya soch rahi hi janab....?

OpenStudy (anonymous):

tm dono ne theek se kiya hai

OpenStudy (paki):

hmmm.... basss sekha karo yaha pe dear... buhat faida hoga... mai khud b seekh raha hun...aur sikha b raha hun..... @shereenkhan

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

okay @paki

OpenStudy (paki):

yup dear...

OpenStudy (anonymous):

:)